OpenStudy (anonymous):

Please help! If the path of a projectile is given by h=-4t^2+256t, determine its maximum height. How would I go about solving this? Options: a. 8192 b. 4096 c. 3072 d. 16,384

OpenStudy (anonymous):

max is a vertex, namely \[-\frac{b}{2a}=-\frac{256}{-4}=16\] then replace x by 16

OpenStudy (anonymous):

are you sure it isn't \[h=-16t^2+256t\]?

OpenStudy (anonymous):

what is h and t ?

OpenStudy (anonymous):

@satellite, Yes I'm sure it isn't ... The problem asks for h=-4t^2+256t. h stands for height but I don't know about t.

OpenStudy (anonymous):

I got C. Can someone verify?

OpenStudy (ash2326):

We have \[h= -4t^2+256 t\] vertex \[\frac{b}{2a}=-\frac{256}{8}=32\] this is the value of t at which height is max \[h= -4t^2+256 t\] t=32 \[h=-4 \times 32^2+256\times 32= 32(256-128)=4096 meters\]

OpenStudy (anonymous):

Oh that makes sense. Thank you for the explanation!